OpenStudy (anonymous):

Write 250,700 in expanded form I got 200+50+700

OpenStudy (anonymous):

I think it might mean in terms of powers of 10, like,\[250,700=2 \times 10^5+5 \times 10^4+0 \times 10^3 + 7 \times 10^2 \]\[+ 0 \times 10^1 + 0 \times 10^0\]

OpenStudy (anonymous):

200 000 + 50 000 +700